# CITYWALK: Enhancing LLM-Based C++ Unit Test Generation via Project-Dependency Awareness and Language-Specific Knowledge

### v1

@@ Abstract.
Unit testing plays a pivotal role in the software development lifecycle, as it ensures code quality. However, writing high-quality unit tests remains a time-consuming task for developers in practice. More recently, the application of large language models (LLMs) in automated unit test generation has demonstrated promising results. Existing approaches primarily focus on interpreted programming languages (e.g., Java), while mature solutions tailored to compiled programming languages like C++ are yet to be explored. The intricate language features of C++, such as pointers, templates, and virtual functions, pose particular challenges for LLMs in generating both executable and high-coverage unit tests. To tackle the aforementioned problems, this paper introduces CITYWALK, a novel LLM-based framework for C++ unit test generation. CITYWALK enhances LLMs by providing a comprehensive understanding of the dependency relationships within the project under test via program analysis. Furthermore, CITYWALK incorporates language-specific knowledge about C++ derived from project documentation and empirical observations, significantly improving the correctness of the LLM-generated unit tests. We implement CITYWALK by employing the widely popular LLM GPT-4o. The experimental results show that CITYWALK outperforms current state-of-the-art approaches on a collection of eight popular C++ projects. Our findings demonstrate the effectiveness of CITYWALK in generating high-quality C++ unit tests.
Unit Test Generation, Large Language Model, Program Analysis, Language-Specific Knowledge, Retrieval-Augmented Generation
@@ 1. Introduction
Although LLM-based approaches have achieved remarkable performance in unit test case generation, the test code generated by LLMs still encounters issues, including compilation errors and assertion failures. Furthermore, current state-of-the-art approaches and tools (Sapozhnikov et al., 2024; Ryan et al., 2024; Chen et al., 2024a; Yuan et al., 2024) primarily concentrate on interpreted programming languages such as Java and Python, with limited mature solutions yet available for generating unit test cases tailored to compiled programming language like C++. Particularly, the intricate language features of C++ present substantial challenges in utilizing LLMs to generate executable unit test cases with high coverage. To underscore the limitations of LLMs in the generation of C++ unit test cases, Figure 1 illustrates three motivating examples that demonstrate typical scenarios in which the most advanced LLM GPT-4o, when provided with the basic prompt  [^1] Following ChatTester (Yuan et al., 2024), the basic prompt comprises the task description for unit test generation, a requirement to understand program’s intent, the source code of the focal method, and the dependency contexts within the focal class file., fails to generate correct unit test cases for the corresponding focal methods (i.e., method under test) in the open-source C++ projects.
@@
  • Limitation 2: Lack of cross-file intended behavior in the corresponding project as guidance for LLMs to generate correct assertions. Considering the example of Focal Method (b), the functionality of the Translate method in the tag class file is to invoke the relevant interface methods provided by the Directives class to translate various YAML tags in accordance with their respective types. In this instance, the variable type is defined within the header file of tag, while the contextual information related to the invoked interface methods resides in the class file of Directives. Given that these cross-file dependency contexts within the yaml-cpp project are not encompassed in the basic prompt, comprehending the functional intent of Translate alone does not sufficiently assist GPT-4o in grasping the usage of Translate. Consequently, Failed Test Case ➋ is unable to configure the correct YAML tag type to align with the target translation pattern. This misalignment results in the actual outcome that deviates from the expected result, thereby leading to an assertion failure in the generated test code.
  • Limitation 3: LLMs struggle to correctly generate complex test code without an understanding of C++ language-specific domain knowledge. In practice, the focal method may need to interact with complex data objects defined in dependent class files. In such cases, developers typically create virtual objects using mocking techniques (Saff and Ernst, 2004; Thomas and Hunt, 2002) to simulate the behavior of real external dependencies. However, to effectively generate such complex test code, LLMs must grasp the principles and applicable scenarios related to mocking (Spadini et al., 2019, 2017). In Focal Method (c), the decode method is required to return a boolean value by verifying whether node is a scalar. As illustrated in the generated Failed Test Case ➌, GPT-4o attempts to simulate a Node class object to mimic the test input. The Error Message indicates that the methods invoked on node are non-virtual functions. As a consequence, GPT-4o generates test code that contains compilation errors, stemming from an insufficient understanding of such complex language feature of C++ mocking.
@@
  • Empowering LLMs with an awareness of project-level dependency relationships through program analysis. The project dependencies employed by CITYWALK are categorized into two types: cross-file data dependencies and configuration dependencies. To conduct cross-file data dependency analysis, CITYWALK initially performs static analysis to identify the header and source code files within the project that exhibit a dependency-chain relationship with the focal class file. Subsequently, CITYWALK utilizes the abstract syntax tree (AST) to extract the relevant cross-file data dependency contexts associated with the focal method from these identified files. Furthermore, our observations indicate that when the project under test necessitates specific versions of the compilation environment or third-party libraries, the probability of encountering compilation errors in the LLM-generated unit tests increases substantially. By analyzing the configuration files of the project, CITYWALK gathers critical information regarding the third-party library usage and the requirements of the compilation environment. These configuration dependencies are explicitly provided to LLMs to reduce the occurrence of compilation errors.
@@
  • Prompting LLMs with step-by-step instructions and language-specific knowledge derived from empirical observations. CITYWALK decomposes the unit test generation task into smaller, more specific subtasks, structured as step-by-step instructions. The goal is to streamline the generation of C++ unit test cases by providing LLMs with detailed procedural guidance. Within CITYWALK, LLMs are utilized to: (1) infer both the intention and dependencies of the focal method; (2) generate a test suite for the focal method by leveraging the provided guidance; and (3) refine the generated test cases using language-specific knowledge. To perform Step (3), we further conduct an empirical analysis of the LLM-generated failed test cases for building language-specific domain knowledge tailored to C++. This knowledge-driven approach assists in guiding LLMs to emulate experienced developers in generating accurate test cases.
We implement a prototype of CITYWALK using GPT-4o and evaluate it on a collected benchmark comprising 1057 focal methods from eight C++ projects. We conduct a comparative analysis of CITYWALK against six state-of-the-art baselines. Our evaluation demonstrates that CITYWALK outperforms all baselines in both compilation success rate and code coverage. We further integrate CITYWALK as a plugin within the open-source code LLMs, with the aim of proving its generalization capabilities and effectiveness in C++ unit test generation. Our contributions can be summarized as follows:
@@
  • We publicly release the artifacts (Zhang, 2024) of CITYWALK on Zenodo to facilitate the reproduction of CITYWALK by the researchers in the software engineering community.
Article Organization. The remainder of this paper is organized as follows: Section 2 introduces in detail the proposed framework. Section 3 and Section 4 provide the experimental setups and results of our research. Section 5 presents case studies and discloses the threats to the validity of our approach. Section 6 describes the related work. Section 7 draws conclusions of our research.
@@ 2.1. Task Formulation
This stage involves navigating through all necessary files in $\mathtt{Repo}$, segmenting the selected files into manageable chunks, and storing the vectorized chunks in the knowledge bases $\mathtt{KB_{code}}$ and $\mathtt{KB_{docs}}$.
@@
This stage involves prompting the LLM with step-by-step instructions to generate test cases $\mathtt{TC\leftarrow LLM(PROMPT)}$, where the content of $\mathtt{PROMPT}$ includes the basic prompt, supplemented with additional guidance from Stage -, and incorporates rule-based methods to fix test cases with syntactic and compilation errors.
@@ 2.2.2. Knowledge Base Construction
Retrieval-augmented generation (RAG) allows LLMs to leverage information from external knowledge bases for reducing hallucinations (Chen et al., 2024b). To enhance the LLM’s understanding of the focal method during unit test generation, CITYWALK constructs knowledge bases using the documentation and source code files from the project repository. CITYWALK first navigates through all representative files within the repository, including project documentation files (e.g., requirement specifications) written in natural language and source code files in programming languages. Next, CITYWALK applies text segmentation strategies to slice the documentation and code files. For instance, markdown files can be segmented into text chunks based on title levels, while source code files are divided into method-level code chunks. Additionally, natural language comments within the code are extracted separately as text chunks. To facilitate efficient semantic retrieval and similarity calculations, CITYWALK utilizes an embedding model BGE (Xiao et al., 2024) for the vectorization of segmented text chunks, storing the resulting vectors in $\mathtt{KB_{docs}}$. The segmented code chunks are directly stored in $\mathtt{KB_{code}}$ for exact match retrieval without vectorization.
@@ 2.3. Project Dependency Extraction
Figure 4. Illustrative Example of Project Dependency Extraction.
@@
2.4. Intention Context Retrieval
To better guide LLMs in understanding the program intention of the focal method, CITYWALK employs RAG to generate effective unit test cases by leveraging the project documentation and source code stored in the knowledge bases described in Section 2.2.2. The natural language documentation provides functionality requirements of the focal method as articulated by developers, while the source code snippets within the project offer LLMs real-world examples of focal method invocation and initialization. This approach mitigates issues related to focal methods that have dependency chains longer than two layers or are complex to initialize. Specifically, CITYWALK employs a hybrid retrieval strategy that conducts knowledge queries in two stages. For $\mathtt{KB_{docs}}$, the process begins by generating a corresponding query statement (as shown in Figure 2) based on the name of the focal method and the class to which it belongs. The embedding model BGE is then utilized to convert this query into a vector representation, capturing its semantic information and enabling alignment with the content of $\mathtt{KB_{docs}}$ in vector space. Using the retrieval algorithms provided by the vector database Faiss (Douze et al., 2024), the similarity between the query vector and the vectors in $\mathtt{KB_{docs}}$ is computed, employing cosine similarity to identify the vectors that are most similar to the query vector. Based on the similarity scores, the top-2 most relevant responses are selected as retrieval results. This setup ensures retrieval accuracy while providing sufficiently detailed guidance information, assisting in the generation of more comprehensive test cases. Preliminary practices indicate that semantic retrieval for code snippet examples often yields imprecise results (Bajracharya et al., 2010; Yan et al., 2020). Thus, CITYWALK optimally employs an exact match retrieval strategy for $\mathtt{KB_{code}}$, using regular expressions to identify examples of method invocations or initializations based on the name of the focal method.
@@ 2.5. Empirical Observation
  • Empirical Setups: (1) Benchmark. We adopt a diverse set of real-world open-source C++ projects collected from GitHub, with the selection criteria and detailed statistics provided in Section 3.2. (2) Subject LLMs. We select two state-of-the-art LLMs to evaluate their capabilities in generating high-quality C++ unit test cases based on our collected benchmark: the open-source LLM DeepSeek-V2.5 (DeepSeek-AI, 2024a, b) and the closed-source LLM GPT-4o (OpenAI, 2023). (3) Prompt Design. As suggested by Yuan et al. (Yuan et al., 2024), we design our prompt by closely following common practices in recent unit test generation research. The prompt consists of a natural language description that explains the task to the LLM, along with the code context, which includes the focal method and other relevant code.
@@
  • Experimental Procedure: We query the selected LLMs using our designed prompt for each focal method within the collected benchmark and consider the test cases generated by the LLMs as the output. The official API of each LLM is used with the configuration set to generate the top-1 chat completion choice and a sampling temperature of 0. The generated test cases are then placed in the test directory of the project under test, where we attempt to compile and execute them for subsequent analysis. We evaluate the correctness of the generated test cases from three perspectives. First, we use the Clang parser as a syntax checker to verify the syntactic correctness of the generated test cases. Next, we measure their compilation and execution correctness by determining whether the test cases could compile successfully and pass execution. In addition, we employ llvm-cov  [^3] https://llvm.org/docs/CommandGuide/llvm-cov.html to collect both line and branch coverage, providing an assessment of the sufficiency of the generated test cases.
  • Error Analysis: To better understand the limitations of existing LLMs in C++ unit test generation, we analyze common error patterns in the failed test cases generated by DeepSeek-V2.5 and GPT-4o. Specifically, we automatically build the error taxonomies based on the error messages encountered during compilation and execution, and then manually group similar errors into high-level categories. As shown in the first two columns of Table 1, we classify the error patterns into three main categories: (A) Compilation Errors, (B) Execution Failures, and (C) Poor Coverage. The LLM-generated test cases exhibit various compilation errors, including import errors, namespace issues, and access errors. Most execution failures stem from incorrectly generated assertions and mocking code. Additionally, we identify poor coverage as a significant issue, primarily due to the LLMs’ inability to ensure coverage of all conditional branches and the incorrect use of assertion types.
  • Solution Guidelines: As listed in the third column of Table 1, we systematically summarize solution guidelines for the corresponding error patterns based on empirical observations of the failed test cases generated by the selected LLMs. As discussed in Section 2.3, Failed Test Case ➌ can be addressed by leveraging additional cross-file data dependencies along with Guideline (B.2). Specifically, we organize Guideline (B.2) as instructions to prompt LLMs in a knowledge-driven manner for tackling the incorrect mocking issue. In doing so, CITYWALK incorporates language-specific knowledge about mocking in C++, including best practices for creating mock objects for virtual functions. In the case of the decode method, CITYWALK provides LLMs with detailed guidance on how to correctly mock non-virtual member functions in C++. By enhancing the LLM’s understanding of these language-specific nuances, CITYWALK ensures the generated mocking code is both syntactically and semantically correct, effectively preventing execution failures caused by improper mocking of non-virtual methods. While these guidelines do not cover all error types, they are instrumental in guiding LLMs to produce high-quality test cases and mitigate common errors in C++ unit test generation.
@@
Category Error Pattern Guideline
(A) Compilation Error Import Error (A.1) Import all necessary dependencies with the correct paths.
@@
Namespace Error (A.4) Use the correct namespace throughout the tests.
Access Error (A.5) Properly handle static members by accessing them using the class name.
@@
(B) Execution Failure Incorrect Assertion (B.1) Choose appropriate assertions for the pointer data type, clearly distinguishing between
address and content comparisons.
Incorrect Mocking (B.2) For mocking (if using gmock), remember that only virtual methods can be mocked.
(C) Poor Coverage Uncovered Branch (C.1) Ensure coverage of true and false branches for each conditional predicate at least once.
Assertions with ASSERT_* (C.2) Utilize non-terminating assertions (e.g., EXPECT_*) to maximize code coverage.

### v2

@@ Abstract.
Unit testing plays a pivotal role in the software development lifecycle, as it ensures code quality. However, writing high-quality unit tests remains a time-consuming task for developers in practice. More recently, the application of large language models (LLMs) in automated unit test generation has demonstrated promising results. Existing approaches primarily focus on interpreted programming languages (e.g., Java), while mature solutions tailored to compiled programming languages like C++ are yet to be explored. The intricate language features of C++, such as pointers, templates, and virtual functions, pose particular challenges for LLMs in generating both executable and high-coverage unit tests. To tackle the aforementioned problems, this paper introduces CITYWALK, a novel LLM-based framework for C++ unit test generation. CITYWALK enhances LLMs by providing a comprehensive understanding of the dependency relationships within the project under test via program analysis. Furthermore, CITYWALK incorporates language-specific knowledge about C++ derived from project documentation and empirical observations, significantly improving the correctness of the LLM-generated unit tests. We implement CITYWALK by employing the widely popular LLM GPT-4o. The experimental results show that CITYWALK outperforms current state-of-the-art approaches on a collection of ten popular C++ projects. Our findings demonstrate the effectiveness of CITYWALK in generating high-quality C++ unit tests.
Unit Test Generation, Large Language Model, Program Dependence Analysis, Language-Specific Knowledge, Retrieval-Augmented Generation
@@ 1. Introduction
Although LLM-based approaches have achieved remarkable performance in unit test case generation, the test code generated by LLMs still encounters issues, including compilation errors and assertion failures. Furthermore, current state-of-the-art approaches and tools (Sapozhnikov et al., 2024; Ryan et al., 2024; Chen et al., 2024a; Yuan et al., 2024) primarily concentrate on interpreted programming languages such as Java and Python, with limited mature solutions yet available for generating unit test cases tailored to compiled programming language like C++. Particularly, the intricate language features of C++ present substantial challenges in utilizing LLMs to generate executable unit test cases with high coverage. To underscore the limitations of LLMs in the generation of C++ unit test cases, Figure 1 illustrates three motivating examples that demonstrate typical scenarios in which the most advanced LLM GPT-4o, when provided with the basic prompt [^1] Following ChatTester (Yuan et al., 2024), the basic prompt comprises the task description for unit test generation, a requirement to understand program’s intent, the source code of the focal method, and the dependency contexts within the focal class file., fails to generate correct unit test cases for the corresponding focal methods (i.e., the methods under test) in the open-source C++ projects.
@@
  • Limitation 2: Lack of cross-file intended behavior in the corresponding project as guidance for LLMs to generate correct assertions. Considering the example of Focal Method (b), the functionality of the Translate method in the tag class file is to invoke the relevant interface methods provided by the Directives class to translate various YAML tags in accordance with their respective types. In this instance, the variable type is defined within the header file of tag, while the contextual information related to the invoked interface methods resides in the class file of Directives. Given that these cross-file dependency contexts within the yaml-cpp project are not encompassed in the basic prompt, comprehending the functional intent of Translate alone does not sufficiently assist GPT-4o in grasping the usage of Translate. Consequently, Failed Test Case ➋ is unable to configure the correct YAML tag type to align with the target translation pattern. This misalignment causes the actual outcome to deviate from the expected result, consequently triggering an assertion failure.
  • Limitation 3: LLMs struggle to correctly generate complex test code without an understanding of C++ language-specific domain knowledge. In practice, the focal method may need to interact with complex data objects defined in dependent class files. In such cases, developers typically create virtual objects using mocking techniques (Saff and Ernst, 2004; Thomas and Hunt, 2002) to simulate the behavior of real external dependencies. However, to effectively generate complex test code, LLMs must grasp the principles and applicable scenarios related to mocking (Spadini et al., 2019, 2017). In Focal Method (c), decode needs to return a boolean value by verifying whether node is a scalar. As illustrated in Failed Test Case ➌, GPT-4o attempts to simulate a Node class object to mimic the test input. The Error Message indicates that the methods invoked on node are non-virtual functions. As a consequence, GPT-4o generates test code that contains compilation errors, stemming from an insufficient understanding of such complex language feature of C++ mocking.
@@
  • Empowering LLMs with an awareness of project-level dependency relationships through program analysis. The project dependencies employed by CITYWALK are categorized into two types: cross-file data dependencies and configuration dependencies. To conduct cross-file data dependency analysis, CITYWALK initially performs static analysis to identify the header and source code files within the project that exhibit a dependency-chain relationship with the focal class file. Subsequently, CITYWALK utilizes the abstract syntax tree (AST) to extract the relevant cross-file data dependency contexts associated with the focal method from these identified files. Furthermore, our observations indicate that when the project under test necessitates specific versions of the compilation environment or third-party libraries, the probability of encountering compilation errors in the LLM-generated unit tests increases substantially. By analyzing project’s configuration files, CITYWALK gathers critical information regarding the third-party library usage and the requirements of the compilation environment. These configuration dependencies are explicitly provided to LLMs to reduce the occurrence of compilation errors.
@@
  • Prompting LLMs with step-by-step instructions and language-specific knowledge derived from empirical observations. CITYWALK decomposes the unit test generation task into smaller, more specific subtasks with structured step-by-step instructions. The goal is to streamline the generation of C++ unit test cases by providing LLMs with detailed procedural guidance. Within CITYWALK, LLMs are utilized to: (1) infer both the intention and dependencies of the focal method; (2) generate an entire test file for the focal method by leveraging the provided guidance; and (3) refine the generated test cases using language-specific knowledge. To perform Step (3), we further conduct an empirical analysis of the LLM-generated failed test cases for building language-specific domain knowledge tailored to C++. This knowledge-driven approach assists in guiding LLMs to emulate experienced developers in generating accurate test cases.
We implement a prototype of CITYWALK using GPT-4o and evaluate it on a collected benchmark comprising 1288 focal methods from ten C++ projects. We conduct a comparative analysis of CITYWALK against seven state-of-the-art baselines. Our evaluation demonstrates that CITYWALK outperforms all baselines in both compilation success rate and code coverage. Our contributions can be summarized as follows:
@@
  • We publicly release the artifacts (Zhang, 2024) of CITYWALK on Zenodo to facilitate the reproduction.
Article Organization. The remainder of this paper is organized as follows: Section 2 introduces in detail the proposed framework. Section 3 and Section 4 provide the experimental setups and results of our research. Section 5 presents multi-perspective discussions and discloses the threats to the validity of our approach. Section 6 describes the related work. Section 7 concludes with research findings and future directions.
@@ 2.1. Task Formulation
This stage involves navigating through all files in $\mathtt{Repo}$, segmenting the selected files into chunks, and storing the vectorized chunks in the knowledge bases $\mathtt{KB_{code}}$ and $\mathtt{KB_{docs}}$.
@@
This stage involves prompting the LLM with step-by-step instructions to generate test cases $\mathtt{TC\leftarrow LLM(PROMPT)}$, where $\mathtt{PROMPT}$ includes the basic prompt supplemented with additional guidance from Stage -, and incorporates a three-phase post-processing approach for error-fixing.
@@ 2.2.2. Knowledge Base Construction
Retrieval-augmented generation (RAG) allows LLMs to leverage information from external knowledge bases for reducing hallucinations (Chen et al., 2024b). To enhance the LLM’s understanding of the focal method during unit test generation, CITYWALK constructs knowledge bases using the documentation and source code from the project repository. First, CITYWALK scans all semantically relevant files within the repository, including project documentation (e.g., requirement specifications) written in natural language and source code files in programming languages. Files that do not contribute semantic knowledge (e.g., configuration files) are excluded from constructing knowledge bases. CITYWALK then applies text segmentation strategies to slice the documentation and code files. For instance, markdown files can be segmented into text chunks based on title levels, while source code files are divided into method-level code chunks. Additionally, natural language comments within the code are extracted separately as text chunks. To facilitate efficient semantic retrieval and similarity calculations, CITYWALK utilizes an embedding model BGE (Xiao et al., 2024) for the vectorization of segmented text chunks, storing the resulting vectors in $\mathtt{KB_{docs}}$. The segmented code chunks are directly stored in $\mathtt{KB_{code}}$ for exact match retrieval without vectorization.
@@ 2.3. Project Dependency Extraction
Figure 4. Illustrative Example of Project Dependency Extraction.
[…]
@@
2.4. Intention Context Retrieval
@@ 2.5. Empirical Observation
  • Empirical Setups: (1) Benchmark. We adopt a diverse set of real-world open-source C++ projects collected from GitHub, with the selection criteria and detailed statistics provided in Section 3.2. (2) Subject LLMs. We select two state-of-the-art LLMs to evaluate their capabilities in generating high-quality C++ unit test cases based on our collected benchmark: the open-source LLM DeepSeek-V3 (DeepSeek-AI, 2024) and the closed-source LLM GPT-4o (OpenAI, 2023). (3) Prompt Design. As suggested by Yuan et al. (Yuan et al., 2024), we design our prompt by closely following common practices in recent unit test generation research. The prompt consists of a natural language description that explains the task to the LLM, along with the code context, which includes the focal method and other relevant contextual information (e.g., the fields and method signatures within the focal class).
@@
  • Experimental Procedure: For each project in the benchmark, we clone its repository from GitHub and extract relevant information to support prompt construction. We query the selected LLMs using our designed prompt for each focal method and consider the test cases generated by the LLMs as the output. We use the official API of each LLM with the configuration set to generate the top-1 chat completion choice and a sampling temperature of 0. The generated test cases are then placed in the test directory of the project, where we attempt to compile and execute them for subsequent analysis. In our experimental design, we evaluate the correctness of the generated test cases from three perspectives. First, we use the Clang parser as a syntax checker to verify the syntactic correctness of the generated test cases. Next, we measure the correctness of compilation and execution by verifying whether the generated test cases compile successfully and run without errors. Error messages produced during compilation and execution are automatically extracted for analysis. To examine the failed test cases generated by the two evaluated LLMs, the first two authors manually classified the errors based on the corresponding compiler-generated messages. This empirical process involved over 1000 failed test cases and required approximately five hours. To ensure consistency and accuracy, the two authors collaboratively reviewed and resolved discrepancies through discussion, reaching consensus on all classification results. In addition, we employ llvm-cov [^3] https://llvm.org/docs/CommandGuide/llvm-cov.html to collect both line and branch coverage, providing an assessment of the sufficiency of the generated test cases.
  • Error Analysis: To better understand the limitations of existing LLMs in C++ unit test generation, we analyze common error patterns in the failed test cases generated by DeepSeek-V3 and GPT-4o, with a particular focus on their impact on compilation correctness. It is important to note that this analysis is limited to compilation errors, which constitute a significant proportion of all errors and represent a critical initial barrier to the successful execution and correctness of generated test cases. Specifically, we automatically categorize each failed test case based on the associated compilation error message. Table 1 presents the distribution of compilation errors observed in the test cases generated by the evaluated LLMs. The “Frequency” column reports the number of occurrences for each error pattern across DeepSeek-V3, GPT-4o, and their combined total. For clarity, only error patterns that appeared more than ten times are included in the table. As shown in Table 1, LLM-generated test cases exhibit a diverse range of compilation error patterns. The most common errors are related to undefined symbols, typically caused by references to unresolved identifiers such as undeclared methods or variables. Two other prevalent error patterns include access violations and type mismatches. Access errors generally result from invalid attempts to reference class members, while type errors stem from incompatible expressions or incorrect type assignments. Additionally, we observe that GPT-4o frequently produces test code with incorrect namespace usage, which is not present in DeepSeek-V3 outputs. This quantitative analysis reveals persistent patterns of compilation errors encountered by LLMs during unit test generation and provides valuable insights that inform the design of our mitigation strategies.
    Table 1. Compilation Error Breakdown for LLM-Generated Test Cases
    Error Pattern Error Description Frequency
    DeepSeek-V3 GPT-4o Total
    Undefined Symbols Error Missing or unresolved identifiers 351 382 733
    Access Error Invalid access to class members 157 147 304
    Type Error Type mismatches in expressions or assignments 140 161 301
    Other Miscellaneous unclassified errors 91 76 167
    Test Setup Error Failure during initialization of tests 51 99 150
    Linker Error Cross-file linkage failure 92 32 124
    Syntax Error Invalid syntax in C++ source files 18 59 77
    Namespace Error Incorrect or missing namespace usage 0 45 45
    Multiple Definition Error Duplicate symbols defined in different files 5 26 31
    Template Error Invalid usage of C++ templates 1 9 10
@@
    Category Guideline
    (A) Compilation Error (A.1) Import all necessary dependencies with the correct paths.
@@
    (A.4) Use the correct namespace throughout the tests.
    (A.5) Properly handle static members by accessing them using the class name.
@@
    (B) Execution Failure (B.1) Choose appropriate assertions for the pointer data type, clearly distinguishing between address and content comparisons.
    (B.2) For mocking (if using gmock), remember that only virtual methods can be mocked.
    (C) Poor Coverage (C.1) Ensure coverage of true and false branches for each conditional predicate at least once.
    (C.2) Utilize non-terminating assertions (e.g., EXPECT_*) to maximize code coverage.
  • Solution Guidelines: As presented in the first column of Table 2, we manually group similar errors into three high-level categories: (A) Compilation Errors, (B) Execution Failures, and (C) Poor Coverage. While compilation errors constitute the majority, the LLM-generated test cases also exhibit execution failures, often caused by incorrect assertions or flawed mocking code. Furthermore, poor coverage emerges as a significant issue, largely attributable to the LLMs’ limited ability to exercise all conditional branches and their misuse of assertion types. As listed in the second column of Table 2, we systematically summarize solution guidelines for the corresponding error patterns based on empirical observations of the failed test cases generated by the selected LLMs. As discussed in Section 2.3, Failed Test Case ➌ can be addressed by leveraging additional cross-file data dependencies along with Guideline (B.2). Specifically, we organize Guideline (B.2) as instructions to prompt LLMs in a knowledge-driven manner for tackling the incorrect mocking issue. In doing so, CITYWALK incorporates language-specific knowledge about mocking in C++, including best practices for creating mock objects for virtual functions. In the case of the decode method, CITYWALK provides LLMs with detailed guidance on how to correctly mock non-virtual member functions in C++. By enhancing the LLM’s understanding of these language-specific nuances, CITYWALK ensures the generated mocking code is both syntactically and semantically correct, effectively preventing execution failures caused by improper mocking of non-virtual methods. While these guidelines do not cover all error types, they are instrumental in guiding LLMs to produce high-quality test cases and mitigate common errors in C++ unit test generation.